timber co bought a machine that cost $ 110,000 on January 1, 2016. The machine had a lifespan of six years and $ 8,000. timber co uses linear depreciation on its machines.
What is a depreciated amortization at the end of 2018 and what are the accumulated amortization at the same time?

Group of answer choices

17,000 depreciated amortization and 51,000 revocation amortization.

17,000 depreciated amortization and 68,000 amortization depreciation.

68,000 depreciated amortization and 17,000 repayment amortization.

17,000 depreciated amortization and 17,000 revocation amortization.
Answer #1

Depreciated amortization

= (110,000-8,000)/6 = 17,000

Accumulated amortization = 17,000 * 3 years

= 51,000

Option A is the answer

Comment if you face any issues